OpenStudy (anonymous):

how do I solve the equation 5x-(2x-1)=2

jhonyy9 (jhonyy9):

can i help to you ?

OpenStudy (anonymous):

yes, I have the awnser just need to know how I did it

jhonyy9 (jhonyy9):

comen i like you see how you will solve it

jhonyy9 (jhonyy9):

firstly you need to eéimine the parantheses

jhonyy9 (jhonyy9):

elimine

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

write me please

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5x-2x-1=2

jhonyy9 (jhonyy9):

not is right

OpenStudy (anonymous):

can you show me how to work this

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

when you have -(-1) what will result to you

OpenStudy (anonymous):

+1

jhonyy9 (jhonyy9):

yes right comen make it now

OpenStudy (anonymous):

5x-2x+1=2

jhonyy9 (jhonyy9):

yes and ...

OpenStudy (anonymous):

or 5x-10x+5=2

jhonyy9 (jhonyy9):

how you have got it this ? what you have wrote 5x-2x+1=2 is right

jhonyy9 (jhonyy9):

cont

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

continue this

OpenStudy (anonymous):

then subtract 1 from bothe sides?

jhonyy9 (jhonyy9):

yes and ?

OpenStudy (anonymous):

5x-2x=2

jhonyy9 (jhonyy9):

why ?

OpenStudy (anonymous):

5x-2x=1 because i subtracted 1 fom both sides

jhonyy9 (jhonyy9):

make what you have wrote before

jhonyy9 (jhonyy9):

yes and 5x-2x=???

OpenStudy (anonymous):

5x-2x=2

jhonyy9 (jhonyy9):

if you subtract 2 from 5 how many will remaind ?

OpenStudy (anonymous):

5x+1

jhonyy9 (jhonyy9):

no there you have wrote 5x-2x=1 so if you subtract 2 from 5 how many will remaind to you ?

OpenStudy (anonymous):

3

jhonyy9 (jhonyy9):

yes and the result ?